can you go through the steps to solve this?
12y-2xy-2x(^2)y
------------
6y-4xy-2x^(2)y

also
20y^(2)-45
-----------
10y^(2)-5y-15

Its asking to reduce in both of the above questions

Answer by flame8855(424)

12y-2xy-2x(^2)y= -2y(x^2+x-6)=-2y(x-2)(x-3)
6y-4xy-2x^2y= -2y(x^2+2x-3)=-2y(x+3)(x-1)
divide both :
12y-2xy-2x(^2)y
------------ = (x-2)/(x-1)
6y-4xy-2x^(2)y

20y^(2)-45 = 5(4y^2-9)=5(2y-3)(2y+3)
10y^(2)-5y-15 = 5(2y^2-y-3)= 5(2y-3)(y+1)
divide both
20y^(2)-45
----------- = (2y+3)/(y+1)
10y^(2)-5y-15